# Verifiable cloud-based variational quantum algorithms

### v2

@@ Abstract
Variational quantum algorithms (VQAs) have shown potential for quantum advantage with noisy intermediate-scale quantum (NISQ) devices for quantum machine learning (QML). However, given the high cost and limited availability of quantum resources, delegating VQAs via cloud networks is a more practical solution for clients with limited quantum capabilities. Recently, Shingu et al.[Physical Review A, 105, 022603 (2022)] proposed a variational secure cloud quantum computing protocol, utilizing ancilla-driven quantum computation (ADQC) for cloud-based VQAs with minimal quantum resource consumption. However, their protocol lacks verifiability, which exposes it to potential malicious behaviors by the server. Additionally, channel loss requires frequent re-delegation as the size of the delegated variational circuit grows, complicating verification due to increased circuit complexity. This paper introduces a new protocol to address these challenges and enhance both verifiability and tolerance to channel loss in cloud-based VQAs.
I Introduction
Quantum computation DiVincenzo (1995) has rapidly transitioned from theoretical speculation to practical application, leveraging the principles of quantum mechanics to tackle problems that are intractable for classical computers. Despite this progress, quantum resources remain scarce and costly, primarily accessible to large corporations. This limitation has spurred efforts to make quantum computation more accessible, especially for clients with limited quantum capabilities.
Blind quantum computation (BQC), a subset of delegated quantum computation (DQC), was first introduced by Childs Childs (2005). It employs quantum one-time padding Deng and Long (2004) in gate-based quantum computation (GBQC), enabling a client to delegate quantum computations while ensuring blindness, i.e., the client’s input, output, and algorithm remain hidden from the server. Building on this concept, Broadbent et al. Broadbent et al. (2009) proposed the universal blind quantum computation (UBQC) protocol, also known as the BFK protocol. This protocol uses brickwork states as resource states within measurement-based quantum computation (MBQC) Briegel et al. (2009) on the server side, requiring the client only to prepare a set of qubits $\left\{\frac{1}{\sqrt{2}}(\ket{0}+e^{i\phi_{k}}\ket{1}\mid\phi_{k}=\frac{k\pi}{4},\,k\in\{0,1,\ldots,7\}\right\}$ while maintaining blindness. This protocol has spurred research in areas such as verification Morimae (2014); Fitzsimons and Kashefi (2017); Barz et al. (2013), the reduction of the client’s quantum capabilities Morimae and Fujii (2013a); Li et al. (2021a); Morimae and Fujii (2013b); Li et al. (2014); Xu and Wang (2021); Zhang (2022); Cao (2023), joint computational tasks Sciarrino et al. (2023), and various applications, including Shor’s algorithm Das and Sanders (2022) and Grover’s algorithm Gustiani and DiVincenzo (2021). Experimental demonstrations of BQC have also been conducted Barz et al. (2012); Huang et al. (2017).
In parallel, variational quantum algorithms (VQAs) Cerezo et al. (2021) have emerged as a promising approach for utilizing noisy intermediate-scale quantum (NISQ) devices to achieve quantum machine learning (QML) Biamonte et al. (2017), demonstrating quantum advantages Khan and Robles-Kelly (2020) in areas such as quantum federated learning (QFL) Chen and Yoo (2021), quantum support vector machines (QSVMs) Rebentrost et al. (2014), and quantum reinforcement learning (QRL) Dong et al. (2008).
Integrating VQAs with BQC provides a promising method for clients with limited quantum capabilities to delegate VQAs to a remote server via cloud networks securely. In previous work, Li et al. Li et al. (2021b) combined the BFK protocol Broadbent et al. (2009) with VQAs to implement delegated QFL. However, this approach requires significant quantum resources, with the server needing to entangle $w\cdot d$ qubits, where $w$ is the number of qubits required by the original NISQ algorithms and $d$ is the depth of the brickwork state Broadbent et al. (2009) in the BFK protocol. Wang et al. Wang et al. (2022) mitigate this by employing qubit reuse Houshmand et al. (2018), lowering the server’s quantum resource consumption to $12w+$ 1 qubits.
Shingu et al.Shingu et al. (2022) further minimized the server’s quantum resource requirements while upholding the principles of BQC by proposing a variational secure cloud quantum computing protocol. This protocol leverages ancilla-driven quantum computation (ADQC)Anders et al. (2010) and the no-signaling principle Morimae and Fujii (2013a) to implement variational quantum algorithms (VQAs) securely. Their approach requires the server to use only $w+1$ qubits per operation. However, this protocol lacks verification and is vulnerable to potential malicious operations by the server. Additionally, it is not robust against channel loss, requiring frequent re-delegation as the size of the delegated variational circuit increases, complicating verification. This paper introduces a new protocol that extends their work by incorporating verification and enhancing tolerance to channel loss while maintaining low quantum resource requirements for the server.
The paper is organized as follows: Section II presents the preliminaries, including ADQC, VQAs, and a review of Shingu et al.’s protocol Shingu et al. (2022). Section III details the proposed protocol. Section IV analyzes the protocol, focusing on verifiability, blindness, correctness, and comparisons with existing protocols. Section V concludes the paper, discussing potential extensions and suggesting directions for future research.
II Preliminaries
@@
II.1 Ancilla-Driven Quantum Computation
Ancilla-driven quantum computation (ADQC) Anders et al. (2010) is a hybrid model that integrates elements of MBQC and GBQC. In ADQC, an ancillary qubit $\ket{+}$ couples with a register qubit to implement a single-qubit gate $J(\phi)$, where $J(\phi)=HR_{z}(\phi)$ and $\phi$ represents the designated rotation angle, or with two register qubits to implement a controlled-Z gate using a fixed operation such as $CZ_{RA}(H_{R}\otimes H_{A})$. Here, $CZ_{RA}$ denotes the controlled-Z gate between the register and ancillary qubits, with $H_{R}$ ($H_{A}$) representing the Hadamard gate for the register (ancillary) qubit. The ancillary qubit is then measured in a specific basis, consistent with MBQC, to ensure determinism, and the measurement outcome determines the evolution of the register qubit(s), as illustrated in Fig. 1 for $J(\phi)$ operator.
Ancillary qubits can be realized as optical photons in optical systems Agarwal (2012) and transmitted to distant locations after being coupled to the register qubits, enabling ADQC to be performed remotely by measuring the ancillary qubits in the basis {|±$\left\{\ket{\pm}^{\phi}\right\}=\left\{\frac{1}{\sqrt{2}}(\ket{0}\pm e^{i\phi}\ket{1})\right\}$.
@@ II.1 Ancilla-Driven Quantum Computation
Figure 1: Circuit for the $J(\phi)$ operator. The prepared qubits are |$\ket{\psi}_{R}$ and |$\ket{+}_{A}$, where the subscripts $R$ and $A$ denote the register qubits and ancillary qubit, respectively. The highlighted section represents the measurement of the ancillary qubit |$\ket{+}_{A}$ at an angle $\phi$ in the $XY$plane, i.e., in the basis {|±$\left\{\ket{\pm}^{\phi}\right\}$. After measurement, the operation $X^{s}HR_{z}(\phi)$ is obtained, where $X$ is the Pauli $X$ operator and $s$ is the measurement result.
II.2 Variational quantum algorithms
Variational quantum algorithms (VQAs) employ parameterized quantum gates, such as $R_{X}$, $R_{Y}$, and $R_{Z}$ Cerezo et al. (2021), to optimize the variational circuit |$\ket{\psi(\vec{\theta})}$, where $\psi(\vec{\theta})$ represents a parametrized wave function.
@@ II.2 Variational quantum algorithms
When using classical data in VQAs, encoding techniques such as amplitude encoding Nakaji et al. (2022), angle encoding Weigold et al. (2020), and hybrid encoding Bhabhatsatam and Smanchat (2023) are employed to transform the data into quantum states. The general structure of the circuit is illustrated in Fig. 2. VQA circuits can be conceptualized as quantum neural networks Jeswal and Chakraverty (2019), where qubits serve as nodes and quantum gates, represented by matrices, correspond to the weights of a neural network, directly influencing the qubit states.
The structure of the variational circuit is expressed as $U(\vec{\theta})=\prod_{i=1}^{n}U_{i}(\vec{\theta^{i}})$, where $\vec{\theta}=\left\{\theta_{1},\theta_{2},\ldots,\theta_{L}\right\}$. Here, each $U_{i}(\vec{\theta^{i}})$ represents the $i$-th variational layer with its corresponding subset of variational parameters $\vec{\theta^{i}}$. The superscripts $\{1,\ldots,n\}$ denote different variational layers, and the subscripts $\{1,\ldots,L\}$ refer to specific variational parameters within the entire set $\vec{\theta}$.
@@
Figure 2: General variational quantum circuit, where all qubits are initialized in the state |$\ket{0}$. The operator $En$ encodes classical data $\vec{x}$ into quantum state $En(\vec{x})\ket{0}^{\otimes w}$, and $U(\vec{\theta})=\prod_{i=1}^{n}U_{i}(\vec{\theta}^{i})$ represents a variational layer with a specific ansatz.
@@
The variational parameters $\vec{\theta}$ in VQAs are iteratively updated and optimized by a classical optimizer, such as Adam Kingma and Ba (2015), utilizing measurement outcomes from the quantum circuit in conjunction with the cost function:
@@
where $E(\vec{\theta})$ represents the expectation value of the output state |$\ket{\psi(\vec{\theta})}$ for a given set of parameters $\vec{\theta}$, and $O$ is the measurement observable.
The parameters $\vec{\theta}=\left\{\theta_{j}\right\}_{j=1}^{L}$ are iteratively updated via gradient descent, such that $\vec{\theta}\leftarrow\vec{\theta}-\eta\nabla C(\vec{\theta})$, where $\eta$ denotes the learning rate. The gradients $\nabla C(\vec{\theta})=\left\{\nabla C(\theta_{j})\right\}_{j=1}^{L}$ are computed using the parameter-shift rule Schuld et al. (2019). Specifically, the gradient $\nabla C(\theta_{j})$ is given by:
$\nabla C(\theta_{j})$ $=\frac{1}{2}\left[C\left(\theta_{j}+\frac{\pi}{2}\right)-C\left(\theta_{j}-\frac{\pi}{2}\right)\right]$ (2)
$=\frac{1}{2}\left[f\left(E\left(\theta_{j}+\frac{\pi}{2}\right)\right)-f\left(E\left(\theta_{j}-\frac{\pi}{2}\right)\right)\right],$
where $\theta_{j}\pm\frac{\pi}{2}=\left\{\theta_{1}\ldots\theta_{j\pm\frac{\pi}{2}}\ldots\theta_{L}\right\}$.
II.3 Review of Shingu et al.’s protocol
The variational secure cloud quantum computing protocol proposed by Shingu et al. employs the circuit depicted in Fig. 1 and necessitates additional gates to execute two-qubit control operations, as shown in Fig. 3.
Figure 3: Circuit for the wo-qubit control gates. The server interacts with the client to obtain 6 $J(\phi)$ operators with parameters $\alpha,\beta,\gamma,$ and $\delta$.
Specifically, a single-qubit gate is implemented using three consecutive $J(\phi)$ operators. In contrast, a two-qubit controlled gate requires the server to perform $CZ$ gates, $H$ gates, and two phase gates: the $S$ gate and the $S^{\dagger}$ gate. The client receives and measures ancillary qubits sent by the server to evolve the register qubits according to ADQC. Subsequently, the server measures the output register qubits and communicates the results to the client.
@@
  2. 2.
    The server measures the output register qubits’ states with $\{\hat{A}^{(c)}_{1},\hat{A}^{(c)}_{2},\ldots,\hat{A}^{(c)}_{K^{(c)}}\}_{c=1}^{G}$ and sends the results to the client via classical communication.
  3. 3.
@@
    The client updates the parameters using its optimizer, such as the parameter-shift rule (Eq. 2), yielding $\vec{\theta}[2]=(\theta_{1}[2],\ldots,\theta_{L}[2])^{T}$ for the next step.
  3. 3.
@@
    The client computes the measurement angles for adopting circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ with $\vec{\theta}[2]$.
  4. 4.
@@
    The client and server reiterate the above steps $(\mathcal{I}-2)$ times with $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ and $\vec{\theta}[j]$. Based on the results from the $j$-th step, classical computations update the client’s parameters to $\vec{\theta}[j+1]$ for $j=2,3,\ldots,\mathcal{I}-1$.
@@
This protocol enables the delegation of VQAs through interaction between the server and the client, requiring only $w$ register qubits and a single ancillary qubit. However, if any of the coupled ancillary qubits are lost during transmission, the circuit needs to be re-delegated, which results in a low tolerance to channel loss. Even the most efficient single-photon detectors in optics, which have shown 99% efficiency in recent studies Kuzanyan et al. (2018); Wenner et al. (2014), are not immune to this issue. For sufficiently large circuits, the probability of re-delegation becomes significant. For example, consider a circuit where each block consists of an average of four single-qubit gates and one two-qubit control gate. A circuit with six blocks would require $4\times 3\times 6+1\times 6\times 6=108$ $J(\phi)$ operators. The probability of needing to re-delegate the circuit is approximately $(1-0.99^{108})\approx 66.22\%$ per delegated $J(\phi)$ operator, increasing exponentially with the number of $J(\phi)$ operators.
@@
Due to channel loss, adding dummy gates for verification becomes impractical, as the circuit would require enough extra trap wires for dummy gates without affecting the wires used for computation. This increases the circuit size and leads to frequent circuit re-delegation. One potential solution to mitigate channel loss is to share Bell pairs between the server and the client, with repeated entanglement generation until successful Morimae and Fujii (2013a). However, in each delegation of the $J(\phi)$ operator, as illustrated in Fig. 1, the ancillary qubit is sent to the client after coupling with the register qubit. While the server can send one half of the Bell pair to the client for measurement, further measurements on the remaining half are required on the server side after coupling with the register qubit. The server cannot complete these measurements in Fig. 1, rendering the use of Bell pairs ineffective in this context.
III The Proposed Protocol

### v3

@@ Abstract
Variational quantum algorithms (VQAs) have shown potential for quantum advantage with noisy intermediate-scale quantum (NISQ) devices for quantum machine learning (QML). However, given the high cost and limited availability of quantum resources, delegating VQAs via cloud networks is a more practical solution for a client with limited quantum capabilities. Recently, Shingu et al. proposed a variational secure cloud quantum computing protocol that leverages ancilla-driven quantum computation (ADQC) to perform cloud-based VQAs with minimal quantum resource consumption. However, their protocol lacks verifiability, which exposes it to potential malicious behaviors by the server. Additionally, channel loss requires frequent re-delegation as the size of the delegated variational circuit grows, complicating verification due to increased circuit complexity. This paper introduces a novel protocol that addresses these challenges by incorporating verifiability and increasing tolerance to channel loss while maintaining low quantum resource consumption for the server and requiring minimal quantum capabilities from the client.
keywords:
@@
Cloud-based variational quantum algorithms , Ancilla-driven quantum computation , Verifiability
journal: Optics and Laser Technology
1 Introduction
Quantum computation [1] has rapidly transitioned from theoretical speculation to practical application, leveraging the principles of quantum mechanics to tackle problems intractable for classical computers. Despite this progress, quantum resources remain scarce and costly, primarily accessible to large corporations. This limitation has spurred efforts to make quantum computation more accessible, especially for a client with limited quantum capabilities.
Blind quantum computation (BQC), a subset of delegated quantum computation (DQC), was first introduced by Childs [2]. It employs quantum one-time padding [3] in gate-based quantum computation (GBQC), enabling a client to delegate quantum computations while ensuring blindness, i.e., the client’s input, output, and algorithm remain hidden from the server. Building on this concept, Broadbent et al. [4] proposed the universal blind quantum computation (UBQC) protocol, also known as the BFK protocol. This protocol uses brickwork states as resource states within measurement-based quantum computation (MBQC) [5] on the server side, requiring the client only to prepare a set of qubits $\left\{\frac{1}{\sqrt{2}}(\ket{0}+exp(i\frac{k\pi}{4})\ket{1})\mid k\in\{0,1,\ldots,7\}\right\}$ while maintaining blindness. This protocol has spurred research in areas such as verification [6, 7, 8], the reduction of the client’s quantum capabilities  [9, 10, 11, 12, 13, 14, 15], joint computational tasks [16], and various applications, including Shor’s algorithm [17] and Grover’s algorithm [18]. Experimental demonstrations of BQC have also been conducted [19, 20].
In parallel, variational quantum algorithms (VQAs)[21] have emerged as a promising framework for leveraging noisy intermediate-scale quantum (NISQ) devices in quantum machine learning (QML)[22]. VQAs have demonstrated potential quantum advantages [23] in domains such as quantum federated learning (QFL)[24], quantum support vector machines (QSVMs)[25], and quantum reinforcement learning (QRL) [26].
Integrating VQAs with BQC provides a promising method for the client with limited quantum capabilities to delegate VQAs to a remote server via cloud networks securely. In previous work, Li et al. [27] combined the BFK protocol [4] with VQAs to implement delegated QFL. However, this approach requires significant quantum resources, with the server needing to entangle $w\cdot d$ qubits, where $w$ is the number of qubits required by the original NISQ algorithms and $d$ is the depth of the brickwork states [4] in the BFK protocol. Wang et al. [28] mitigated this by employing qubit reuse [29], lowering the server’s quantum resource consumption to $12w+$ 1 qubits.
Shingu et al.[30] further minimized the server’s quantum resource consumption while upholding the principles of BQC by proposing a variational secure cloud quantum computing protocol. This protocol leverages ancilla-driven quantum computation (ADQC)[31] and the no-signaling principle [9] to implement variational quantum algorithms (VQAs) securely. Their approach requires the server to use only $w+1$ qubits per operation. However, this protocol lacks verification and is vulnerable to potential malicious operations by the server. Additionally, it is not robust against channel loss, requiring frequent re-delegation as the size of the delegated variational circuit increases, complicating verification. This paper introduces a new protocol that extends their work by incorporating verifiability and increasing tolerance to channel loss while maintaining low quantum resource consumption for the server and requiring minimal quantum capabilities from the client.
The paper is organized as follows: Section 2 covers the preliminaries, including ADQC, VQAs, and a review of Shingu et al.’s protocol [30]. Section 3 describes the proposed protocol. Section 4 provides an analysis of the protocol, with a focus on verifiability, blindness, correctness, and comparisons with existing protocols. Finally, Section 5 concludes the paper by discussing potential extensions and suggesting directions for future research.
@@ 1 Introduction
Figure 1: Circuit for the $J(\phi)$ operator: The prepared qubits are |$\ket{\psi}_{R}$ and |$\ket{+}_{A}$, where the subscripts $R$ and $A$ denote the register qubits and ancillary qubit, respectively. The highlighted section represents the measurement of the ancillary qubit |$\ket{+}_{A}$ in the basis $\left\{\frac{1}{\sqrt{2}}(\ket{0}\pm exp(-i\phi)\ket{1})\right\}$. After measurement, the operation $X^{s}HR_{Z}(\phi)$ is obtained, where $X$ is the Pauli $X$ operator and $s$ is the measurement result.
2 Preliminaries
2.1 Ancilla-Driven Quantum Computation
Ancilla-driven quantum computation (ADQC) [31] is a hybrid model that integrates elements of MBQC and GBQC. In ADQC, an ancillary qubit $\ket{+}$ is coupled with a register qubit to implement a single-qubit operation $J(\phi)$:
$J(\phi)=HR_{Z}(\phi),$ (1)
where $\phi$ represents the designated rotation angle. Alternatively, it can be coupled with two register qubits to implement a controlled-Z gate using the fixed coupling operation $(H_{R}\otimes H_{A})CZ_{RA}$. Here, $CZ_{RA}$ denotes the controlled-Z gate between the register and ancillary qubits, with $H_{R}$ ($H_{A}$) representing the Hadamard gate applied to the register (ancillary) qubit. The ancillary qubit is then measured in a specific basis, consistent with MBQC, to ensure determinism, and the measurement outcome determines the evolution of the register qubit(s), as illustrated in Fig. 1 for $J(\phi)$ operator.
Ancillary qubits can be realized as optical photons in optical systems [32] and transmitted to distant locations after being coupled to the register qubits, enabling ADQC to be performed remotely by measuring the ancillary qubits in the basis:
$M(-\phi)=\left\{\frac{1}{\sqrt{2}}(\ket{0}\pm exp(-i\phi)\ket{1})\right\}.$ (2)
2.2 Variational quantum algorithms
Variational quantum algorithms (VQAs) utilize parameterized quantum gates, such as $R_{X}$, $R_{Y}$, and $R_{Z}$ [21], which rotate qubits around the x, y, and z axes of the Bloch sphere, respectively. The rotation angles in the variational circuit |$\ket{\psi(\vec{\theta})}$ are optimized, where |$\ket{\psi(\vec{\theta})}$ denotes a parameterized quantum state.
@@ 2.2 Variational quantum algorithms
When incorporating classical data into VQAs, encoding techniques such as amplitude encoding [33] are used to transform the data into quantum states. The general structure of the circuit is illustrated in Fig. 2. VQA circuits can be conceptualized as quantum neural networks [34], where qubits serve as nodes, and quantum gates, represented by matrices, correspond to the weights of a neural network, directly influencing the qubit states.
@@
Figure 2: General variational quantum circuit: All qubits are initialized in the state |$\ket{0}$. The operator $En$ encodes classical data $\vec{x}$ into the quantum state $En(\vec{x})\ket{0}^{\otimes w}$. The unitary operator $U(\vec{\theta})=\prod_{i=1}^{n}U_{i}(\vec{\theta}^{i})$ represents the variational layers, forming a specifically designed ansatz, where each $U_{i}(\vec{\theta^{i}})$ corresponds to the $i$-th layer of the $n$ variational layers.
@@
The variational parameters $\vec{\theta}=\left\{\theta_{1},\theta_{2},\ldots,\theta_{L}\right\}$, where $L$ denotes the total number of parameters, are iteratively updated and optimized by a classical optimizer, such as Adam [35], using measurement outcomes from the quantum circuit along with the cost function:
@@
where $E(\vec{\theta})$ denotes the expectation value of the output state |$\ket{\psi(\vec{\theta})}$ for a given set of parameters $\vec{\theta}$, and $O$ is the measurement observable. The function $f$ represents the objective function, such as mean squared error (MSE) or cross-entropy.
The parameters $\vec{\theta}=\left\{\theta_{j}\right\}_{j=1}^{L}$ are iteratively updated via gradient descent, such that $\vec{\theta}\leftarrow\vec{\theta}-\eta\nabla C(\vec{\theta})$, where $\eta$ denotes the learning rate. The gradients $\nabla C(\vec{\theta})=\left\{\frac{\partial C(\vec{\theta})}{\partial\theta_{j}}\right\}_{j=1}^{L}$ are computed using the parameter-shift rule [36]. Specifically, the gradient $\frac{\partial C(\vec{\theta})}{\partial\theta_{j}}$ is given by:
$\frac{\partial C(\vec{\theta})}{\partial\theta_{j}}$ $=\frac{\partial f(E(\vec{\theta}))}{\partial E(\vec{\theta})}\cdot\frac{\partial E(\vec{\theta})}{\partial\theta_{j}}$ (4)
$=\frac{1}{2}\frac{\partial f(E(\vec{\theta}))}{\partial E(\vec{\theta})}\cdot\left[E\left(\theta_{j+}\right)-E\left(\theta_{j-}\right)\right],$
where $\theta_{j\pm}=\left\{\theta_{1},\ldots,\theta_{j}\pm\frac{\pi}{2},\ldots,\theta_{L}\right\}$.
2.3 Review of Shingu et al.’s Protocol
The variational secure cloud quantum computing protocol proposed by Shingu et al. employs the circuit depicted in Fig. 1 to perform single-qubit gates and necessitates additional gates to execute two-qubit controlled gates, as shown in Fig. 3.
Figure 3: Circuit for the two-qubit controlled gates: The server interacts with the client to obtain 6 $J(\phi)$ operators with parameters $\alpha$, $\beta$, $\gamma$, and $\delta$.
Specifically, a single-qubit gate $U$ is implemented using three consecutive $J(\phi)$ operators with $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$, following Euler’s rotation theorem [37]. Here, $U=J(\phi_{1})J(\phi_{2})J(\phi_{3})$ and $HU=R_{Z}(\phi_{1})R_{X}(\phi_{2})R_{Z}(\phi_{3})$, where $\phi_{1}$, $\phi_{2}$, and $\phi_{3}$ are Euler angles. In contrast, a two-qubit controlled gate $CU$ is realized by decomposing it into multiple $J(\phi)$ operators to be delegated, along with additional $CZ$ gates, Hadamard ($H$) gates, the $S$ gate, and its conjugate transpose, the $S^{\dagger}$ gate, which are directly performed by the server, as shown in Fig. 3. The client receives and measures ancillary qubits sent by the server to evolve the register qubits according to ADQC. Subsequently, the server measures the output register qubits and sends the results to the client.
@@ 1. A2-1:
  2. A2-2:
    The server measures the output register qubits’ states with the measurement observables $\{\hat{A}^{(c)}_{1},\hat{A}^{(c)}_{2},\ldots,\hat{A}^{(c)}_{K^{(c)}}\}_{c=1}^{G}$ and sends the results to the client via classical communication.
  3. A2-3:
    The client compensates for the Pauli byproduct effect.
  4. A2-4:
@@ 2. A3-2:
    The client updates the parameters using its optimizer according to Eq. 4, resulting in $\vec{\theta}[2]=(\theta_{1}[2],\ldots,\theta_{L}[2])^{T}$ for the next iteration.
  3. A3-3:
@@
    The client computes the measurement angles for the circuits $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ using the parameters $\vec{\theta}[2]$.
  4. A3-4:
@@
    The client and server reiterate the above steps $(\mathcal{I}-2)$ times with $\{U^{(c)}_{\rm{AN}}\}_{c=1}^{G}$ and $\vec{\theta}[j]$. Based on the results from the $j$-th step, the client’s optimizer updates the parameters to $\vec{\theta}[j+1]$ for $j=2,3,\ldots,\mathcal{I}-1$.
@@
This protocol enables the delegation of VQAs through interaction between the server and the client, requiring only $w$ register qubits and a single ancillary qubit. However, if any of the coupled ancillary qubits are lost during transmission, the circuit needs to be re-delegated, which results in a low tolerance to channel loss. Even the most efficient single-photon detectors in optics, which have shown 99% efficiency in recent studies [38, 39], are not immune to this issue. For sufficiently large circuits, the probability of re-delegation becomes significant. For example, consider a circuit where each block consists of an average of four single-qubit gates and one two-qubit control gate. A circuit with six blocks would require $4\times 3\times 6+1\times 6\times 6=108$ $J(\phi)$ operators. The probability of needing to re-delegate the circuit is approximately $(1-0.99^{108})\approx 66.22\%$ per delegated $J(\phi)$ operator, increasing exponentially with the number of $J(\phi)$ operators.
@@
Due to channel loss, adding dummy gates for verification becomes impractical, as the circuit would require enough extra trap wires for dummy gates without affecting the wires used for computation. This increases the circuit size and leads to frequent circuit re-delegation. One potential solution to mitigate channel loss is to share Bell pairs between the server and the client, with repeated entanglement generation until successful [9]. However, during each delegation of the $J(\phi)$ operator, as illustrated in Fig. 1, the ancillary qubit is sent to the client for measurement after being coupled with the register qubit. Although the server can send one half of the Bell pair to the client for measurement, additional measurements on the remaining half are required on the server side after being coupled with the register qubit. However, the server cannot perform these measurements in the context of Fig. 1, where measurement on the ancillary qubit is performed by the client, rendering the use of Bell pairs ineffective in this context.
3 The Proposed Protocol